OpenStudy (artfis):

I'm confused. I keep getting a weird fraction. Help on what I'm doing wrong or what I should do please? {y = 5/2x} {x + 3y = 3}

hero (hero):

If y equals a fraction of x, chances are, the result will also be a fraction.

OpenStudy (anonymous):

{y = 5/2x} {x + 3y = 3} x+15/2x=3 * 2x 2x^2+15=6x 2x^2-6x+15=0 then D=36-4.15.2 < 0 hence your equations don't have real roots, only complex ones if asked for real solutions just write that the system has none

hero (hero):

Here's an approach: y = (5/2)x y = (5x)/2 2y = 5x x + 3y = 3 5x + 15y = 15 2y + 15y = 15 17y = 15 `y = 15/17` 2(15/17) = 5x 30/17 = 5x `x = 6/17`

OpenStudy (anonymous):

and here's what's wrong with your approach you take it that its y=(5x)/2 and i think it's 5/(2x)

OpenStudy (artfis):

Makes more sense. How did you get the 2y + 15y = 15 after the 5x + 15y = 15?

hero (hero):

Because 2y = 5x, so you just replace 5x with 2y in the second equation.

hero (hero):

@mivanov, you guessed incorrectly.

OpenStudy (artfis):

@Hero Thank you very much.